OpenStudy (anonymous):

help pls:)) Eight less than the half the number of students in the freshman class have their own calculators? if 258 calculators are owned by freshmen, how many freshmen are there??

OpenStudy (vijay):

1

OpenStudy (anonymous):

can you show me how??

OpenStudy (anonymous):

258

OpenStudy (vijay):

u will get x in minus :P

OpenStudy (anonymous):

just understand the question well and youl get the answer

OpenStudy (anonymous):

pls. show me your solutions

OpenStudy (saifoo.khan):

(258 * 2 ) - 8 516 - 8 508 students.

OpenStudy (saifoo.khan):

or freshman.

OpenStudy (anonymous):

2(258+8) ??

OpenStudy (anonymous):

i did not do any solutions..

OpenStudy (anonymous):

is that the ans?

OpenStudy (anonymous):

the answer is 258

OpenStudy (saifoo.khan):

HOW?

OpenStudy (anonymous):

by using the .. let x.... etc.

OpenStudy (saifoo.khan):

idk. He should show working.

OpenStudy (anonymous):

the question is how many FRESHMEN are there..in the first sentence it say that 8 less than the half of FRESHMAN...

OpenStudy (saifoo.khan):

(258 * 2 ) - 8 OR (258 * 2 ) + 8 516 - 8 OR 516 + 8 508 freshman. OR 524 freshman

OpenStudy (saifoo.khan):

One of these two.

OpenStudy (anonymous):

nothing lol..

OpenStudy (anonymous):

i still don't get the answer...

OpenStudy (anonymous):

the answer is 258 freshmen..

OpenStudy (anonymous):

yeah but why? illogical...

OpenStudy (anonymous):

yeah its logical

OpenStudy (anonymous):

just understand the problem...

OpenStudy (saifoo.khan):

IT can't be 258, you have to double that, then add/subtract 8.

OpenStudy (anonymous):

the first line says eight less than HALF THE NUMBER OF STUDENTS in the freshman class

OpenStudy (saifoo.khan):

RIght, and we are going in Reverse.

OpenStudy (anonymous):

yeah its FRESHMAN..the middle line say 258 where owned by FRESHMEN...then the question is how many FRESHMEN in there...

OpenStudy (anonymous):

understand that well...

OpenStudy (anonymous):

sorry but than it would be 8 less than 1/2 freshman which is -7 1/2 freshman...

OpenStudy (saifoo.khan):

I agree with Jay.

OpenStudy (saifoo.khan):

Vinxc is mixing up.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

lol..so you guys did not get my idea...

OpenStudy (anonymous):

i definitely did not...

OpenStudy (saifoo.khan):

Same here.

OpenStudy (anonymous):

its up to you guys...im done with this..

OpenStudy (anonymous):

Let x be all the students and y those who have calculators. \[y=\frac{x}{2}-8\]\[y=258\]\[\frac{x}{2}-8=258\]\[x=125\]The answer is that there are 125 students in the class.

OpenStudy (saifoo.khan):

this means i was right in the first attempt. :D

OpenStudy (anonymous):

but then it would be x/2 - 8 = 258 --> x/2 = 258 + 8 --> x = 2(258+8)

OpenStudy (saifoo.khan):

Agree with Jay.

OpenStudy (anonymous):

Yes!

OpenStudy (saifoo.khan):

Let's solve 125 in your equation: 2(125 + 8) 266 Wrong answer.

OpenStudy (anonymous):

u have to add 8 on the right side, not subtract

OpenStudy (anonymous):

Number of freshmen who have their calculator 1/2x -8 The number of students who have there calculator 258 x is the number of students x is the number of students It wants to know how many freshmen are there 1/2x -8 = 258

OpenStudy (anonymous):

agree with doodle

OpenStudy (anonymous):

I'm sorry let me correct this: \[...\frac{x}{2}-8=258\]\[\frac{x}{2}=266\]\[x=2\times266\]\[x=532\]There are 532 students in the class.

OpenStudy (anonymous):

i got the same answer... vinxc u agree with doodle

OpenStudy (anonymous):

as in same ans as doodle, not kmousou

OpenStudy (anonymous):

i mean same as kmousou!!!

OpenStudy (anonymous):

yeah i agree with doodle

OpenStudy (anonymous):

what is the right one?

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

the solution of doodle..

OpenStudy (anonymous):

so 258 is the answer??

OpenStudy (anonymous):

258 freshmen

OpenStudy (anonymous):

First sentence Eight less than the half the number of students in the freshman class have their own calculators This tells us this many have their Own calculator 1/2x -8 Below also tell us 258 have their OWN calculator if 258 calculators are owned by freshmen, how many freshmen are there?? so if 1/2x - 8 own their calculators and 258 own their calculator the x is the number of students because where it says half the number of students = 1/2x 1/2x - 8 = 258 just solve the equation now

OpenStudy (anonymous):

nice doodle

OpenStudy (anonymous):

solve the equation 1/2x - 8 = 258 2(1/2x - 8) = 2(258) 1x - 16 = 516 1x = 516 + 16 1x/1 = 532/1 x = 532 students

OpenStudy (anonymous):

We can check this by \[ \frac{1}{2}x * 532 -8 = 258 \] calculators

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (anonymous):

258

OpenStudy (anonymous):

the final answer is there was 532 students/freshmen

OpenStudy (anonymous):

haha..just get the 2 answers so that you can be confuse..ROTF..

OpenStudy (anonymous):

i have a question: Isn't it 1x is consider as x right?

OpenStudy (anonymous):

it is 532... don't trust others

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Read what I have posted and how I have broke it down and you will see that 532 is correct

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

Kyragay Yes 1x = x